.   Suppose the demand function facing a monopolist is given as p=50-q and the cost function is given as C=10q.   Determine the total revenue function for the monopolist, and graph the relationship between total revenue, total cost and output. In your answer determine the profit maximizing level of output that would be chosen by the monopolist, and discuss mathematically how you know it is a profit maximum.
Answer #1

Given, p= 50-9 50-9 , C(q) = 109 Rla) - þq |R(q) = 50q-q2 l Graph between R(q) and c(q). C(q) = 109 (40,400) 100,0) R(9)=509-